Mathematics 242 – Principles of Analysis Definitions Quiz 2 – February 8, 2013

Your Name:

A) (5) What does the Well Ordering Property of N say?

B) (5) Let f(x) = |x − 3|. What is lub{f(x) | x ∈ (−2, 4)}?